# Preserving Data Secrecy in Decentralized Inter-organizational Process Mining

### 05 Jan 2024 v2

@@ 5.2 The CONFINE protocol
We orchestrate the interaction of the components in CONFINE via a protocol. We separate it in four subsequent stages, namely (i) initialization, (ii) remote attestation, (iii) data transmission, and (iv) computation. These stages are depicted in Figs. 4(a), 4(b), 4(c) and 6, respectively. They are mainly enacted by a Miner Node and Provisioner Nodes. We assume their communication channel is reliable [10] and secure [21]. In the following, we describe each phase in detail.
@@
Remote attestation. The remote attestation serves the purpose of establishing trust between miners and provisioners in the context of fulfilling data requests. This phase adheres to the overarching principles outlined in the RATS RFC standard [9] serving as the foundation for several TEE attestation schemes (e.g., Intel EPID,111sgx101.gitbook.io/sgx101/sgx-bootstrap/attestation. Accessed: January 5, 2024. and AMD SEV-SNP222amd.com/en/processors/amd-secure-encrypted-virtualization. Accessed: January 5, 2024.). Remote attestation has a dual objective: (i) to furnish provisioners with compelling evidence that the data request for an event log originates from a Trusted App running within a TEE; (ii) to confirm the specific nature of the Trusted App as an authentic Secure Miner software entity. This phase is triggered when the Log Requester sends a new case request to the Log Server, specifying: (i) the segment size (henceforth, seg_size), and (ii) the set of the requested case references. Both parameters will be used in the subsequent data transmission phase. Each of the Log Servers commences the verification process by requesting the necessary information from the Log Receiver to conduct the attestation (2). Subsequently, the Log Receiver generates the attestation report containing the so-called measurement of the Trusted App, which is defined as the hash value of the combination of its source code and data. Once this report is signed using the attestation private key associated with the TEE’s hardware of the Miner Node, it is transmitted by the Log Receiver to the Log Servers alongside the attestation public key of the Miner Node (3). The Log Servers authenticate the miner using the public key and decrypt the report (4). In this last step, the Log Servers undertake a comparison procedure in which they juxtapose the measurement found within the decrypted report against a predefined reference value associated with the source code of the Secure Miner. If the decrypted measurement matches the predefined value, the Miner Node gains trust from the provisioner.
@@
Data transmission. Once the trusted nature of the Trusted App is verified, the Log Servers proceed with the transmission of their cases. To accomplish this, each Log Server retrieves the event log from the PAIS (steps 1 and 2, in Fig. 4(c)), and filters it according to the case reference set specified by the miner. Given the constrained workload capacity of the TEE, it is imperative for Log Servers to partition the filtered event log into distinct segments. Consequently, each Log Server generates log segments comprising a variable count of entire cases (3). The cumulative size of these segments is governed by the threshold parameter specified by the miner in the initial request (step 1 of the remote attestation phase, Fig. 4(b)). As an illustrative example from our motivating scenario, the Log Server of the hospital may structure the segmentation such that and reside within the same segment, whereas the specialized clinic might have and in separate segments. Subsequently, the Log Servers transmit their encrypted segments to the Log Receiver of the Trusted App (4). The Log Receiver, in turn, collects the responses in a queue, processing them one at a time. After decrypting the processed segment (5), the Log Receiver forwards the cases contained in the segment to the Log Manager (6). To reconstruct the process instance, cases belonging to the same process instance must be merged by the Log Manager resulting in a single trace (e.g., for Alice case) comprehensive of all the events in the partial traces (e.g., , and for Alice case). During this operation, the Log Manager applies a specific merging schema (i.e., a rule specifying the attributes that link two cases during the merge) as stated in [11]. In our illustrative scenario, the merging schema to combine the cases of Alice is contingent upon the linkage established through their case identifier (i.e., 312). We underline that our proposed solution facilitates the incorporation of diverse merging schemas encompassing distinct trace attributes. The outcomes arising from merging the cases within the processed segment are securely stored by the Log Manager in the TEE.
@@
Computation. The Trusted App requires all the provisioners to have delivered cases referring to the same process instances. For example, when the hospital and the other organizations have all delivered their information concerning case 312 to the Trusted App, the process instance associated with Alice becomes eligible for computation. Upon meeting this condition, the Log Manager forwards the case earmarked for computation to the Log Elaborator (step 1 in Fig. 6). Subsequently, the Log Elaborator proceeds to input the merged case into the process mining algorithm (2). Ultimately, the outcome of the computation is relayed by the Log Elaborator from the TEE to the TEE Interface running atop the Operating System of the Miner Node (3). The CONFINE protocol does not impose restrictions on the post-computational handling of results. In our motivating scenario, the University and the National Institute of Statistics, serving as miners, disseminate the outcomes of computations, generating analyses that benefit the provisioners (though the original data are never revealed in clear). Furthermore, our protocol enables the potential for provisioners to have their proprietary Secure Miner, allowing them autonomous control over the computed results.
@@ 5.3 Implementation
We implemented the Secure Miner component as an Intel SGX333sgx101.gitbook.io/sgx101/. Accessed: January 5, 2024. trusted application, encoded in Go through the EGo framework.444docs.edgeless.systems/ego. Accessed: January 5, 2024. We resort to a TLS [30] communication channel between miners and provisioners over the HTTP web protocol to secure the information exchange. To demonstrate the effectiveness of our framework, we re-implemented and integrated the HeuristicsMiner discovery algorithm [31] within the Trusted Application. Our implementation of CONFINE, including the HeuristicsMiner in Go, is openly accessible at the following URL: github.com/Process-in-Chains/CONFINE/.
@@ 6 Evaluation
In this section, we evaluate our approach through the testing of our tool implementation. We begin with a convergence analysis to demonstrate the correctness of the collaborative data exchange process. Subsequently, we gauge the memory usage with synthetic and real-life event logs, to observe the trend during the enactment of our protocol and assess scalability. We recall that we focus on memory utilization since the availability of space in the dedicated areas is limited as we discussed in Sect. 5.1. We discuss our experimental results in the following. For the sake of reproducibility, we make available all the testbeds and results in our public code repository (linked above).
Output convergence. To experimentally validate the correctness of our approach in the transmission and computation phases (see Sect. 5), we run a convergence test. To this end, we created a synthetic event log consisting of cases of events on average (see Table 2) by simulating the inter-organizational process of our motivating scenario (see Fig. 1)555We generated the event log through BIMP (https://bimp.cs.ut.ee/). We filtered the generated log by keeping the sole events that report on the completion of activities, and removing the start and end events of the pharmaceutical company and specialized clinic’s sub-processes. and we partitioned it in three sub-logs (one per involved organization), an excerpt of which is listed in Table 1. We run the stand-alone HeuristicsMiner on the former, and processed the latter through our CONFINE toolchain. As expected, the results converge and are depicted in Fig. 7 in the form of a workflow net [1]. For clarity, we have colored activities recorded by the organizations following the scheme of Table 2 (black for the hospital, blue for the pharmaceutical company, and green for the specialized clinic).
@@
Memory usage. Figures 7(a) and 7(b) display plots corresponding to the runtime memory utilization of our CONFINE implementation (in MegaBytes). Differently from Fig. 7(b), Fig. 7(a) excludes the computation stage by leaving the HeuristicsMiner inactive so as to isolate the execution from the mining-specific operations. The dashed lines mark the starting points for the remote attestation, the data transmission and the computation stages. We held the seg_size constant at KiloBytes. We observe that the data transmission stage reaches the highest peak of memory utilization, which is then partially freed by the subsequent computation stage, steadily occupying memory space at a lower level. To verify whether this phenomenon is due to the synthetic nature of our simulation-based event log, we also gauge the runtime memory usage of two public real-world event logs too (Sepsis [24] and BPIC 2013 [29]). The characteristics of the event logs are summarized in Table 2. Since those are intra-organizational event logs, we split the contents to mimic an inter-organizational context. In particular, we separated the Sepsis event log based on the distinction between normal-care and intensive-care paths, as if they were conducted by two distinct organizations. Similarly, we processed the BPIC 2013 event log to sort it out into the three departments of the Volvo IT incident management system. Figure 7(c) depicts the results. We observe that the processing of the BPIC 2013 event log demands more memory, particularly during the initial stages, probably owing to its larger size. Conversely, the Sepsis event log turns out to entail the least expensive run. To verify whether these trends are affected by the dimension of the exchanged data segments, we conducted an additional test to examine the trend of memory usage as the seg_size varies with all the aforementioned event logs. Notably, the polylines displayed in Fig. 7(d) indicate a linear increment of memory occupation until a breakpoint is reached. After that, the memory in use is steady. These points, marked by vertical dashed lines, correspond to the seg_size value that allows the provider’s segments to be contained in a single data segment.
@@
Scalability. In this subsection, we examine the scalability of the Secure Miner, focusing on its capacity to efficiently manage an increasing workload in the presence of limited memory resources. We implemented three distinct test configurations gauging runtime memory usage as variations of our motivating scenario log. In particular, we considered (I) the maximum number of events per case, (II) the number of cases , and (III) the number of provisioning organizations as independent integer variables. To conduct the test on the maximum number of events, we added a loop back from the final to the initial activity of the process model, progressively increasing the number of iterations at a step of , resulting in events. Concerning the test on the number of cases, we simulated additional process instances so that having . Finally, for the assessment of the number of organizations, the test necessitated the distribution of the process model activities’ into a variable number of pools, each representing a different organization (). We parameterized the above configurations with three segment sizes (in KiloBytes): for tests I and II, and for test III (the range is reduced without loss of generality to compensate the partitioning of activities into multiple organizations). To facilitate a more rigorous interpretation of the output trends across varying seg_size configurations, we employ two well-known statistical measures. As a primary measure of goodness-of-fit, we employ the coefficient of determination [6], which assesses the degree to which the observed data adheres to the linear () and logarithmic () regressions derived from curve fitting approximations. To further delve into the analysis of trends with a high , we consider the slope of the approximated linear regression [5].
@@
Table 9(d) lists the measurements we obtained. We describe them to elucidate the observed patterns. Figure 8(a) depicts the results of test I, focusing on the increase of memory utilization when the number of events in the event logs grows. We observe that the memory usage for seg_size and (depicted by green and lilac lines, respectively) are quite similar, whereas the setting with seg_size (blue line) exhibits significantly higher memory usage. For the settings with seg_size and , approaches , signifying an almost perfect approximation of the linear relation, against lower values. In these test settings, is very low yet higher than , thus indicating that memory usage is likely to continue increasing as the number of max events grows. The configuration with seg_size yields a higher value, thus suggesting a logarithmic trend, hence a greater likelihood of stabilizing memory usage growth rate as the number of maximum events increases. In Fig. 8(b), we present the results of test II, assessing the impact of the number of cases on the memory consumption. As expected, the configurations with seg_size set to and exhibit a trend of lower memory usage than settings with seg_size . The score of the trends with seg_size and indicate a strong linear relationship between the dependent and independent variables compared to the trend with seg_size , which is better described by a logarithmic regression (). For the latter, the value is higher than the corresponding thus suggesting that the logarithmic approximation is better suited to describe the trend. Differently from test I, the score associated with the linear approximations of the trends with seg_size and approaches , indicating that the growth rate of memory usage as the number of cases increases is negligible. In Fig. 8(c), we present the results of test III, on the relation between the number of organizations and the memory usage. The chart shows that memory usage trends increase as provisioning organizations increase for all three segment sizes. The values for the three seg_sizes are very high, indicating a strong positive linear correlation. The test with seg_size exhibits the slowest growth rate, as corroborated by the lowest result (). For the configuration with seg_size , the memory usage increases slightly faster (). With seg_size , the overall memory usage increases significantly faster than the previous configurations (). We derive from these findings that the Secure Miner may encounter scalability issues when handling settings with a large number of provisioning organizations. Further investigation is warranted to determine the precise cause of this behavior and identify potential mitigation strategies.
In the next section, we conclude our work and outline future research directions based upon our current findings and the limitations of our approach.
@@ 7 Conclusion and Future Work
Confidentiality is paramount in inter-organizational process mining due to the transmission of sensitive data across organizational boundaries. Our research investigates a decentralized secrecy-preserving approach that enables organizations to employ process mining techniques with event logs from multiple organizations while ensuring the protection of privacy and confidentiality. Our solution offers a number of directions to walk along for improvement. We operate under the assumption of fair conduct by data provisioners and do not account for the presence of injected or maliciously manipulated event logs. In addition, we assume that miners and provisioners exchange messages in reliable communication channels where no loss or bit corruption occurs. Our approach relies on certain assumptions about event log data, including the existence of a universal clock for event timestamps, which may not be realistic in situations where organizations are not perfectly synchronized. We aim at enhancing our approach to make it robust to the relaxation of these constraints. Our future work encompasses the integration of usage control policies that specify rules on event logs’ utilization. We plan to design policy enforcement and monitoring mechanisms to achieve this goal following the principles already addressed in [7, 8]. Our solution embraces process mining techniques in a general way. However, we believe the presented approach is compatible with declarative model representations [14]. Therefore, trusted applications could compute and store the entire set of rules representing a business process, and users may interact with them via trusted queries. Finally, in our implementation, we have focused on process discovery tasks. Nevertheless, our approach has the potential to seamlessly cover a wider array of process mining functionalities such as conformance checking, and performance analysis techniques. Implementing them and showing their integrability with our approach paves the path for future research endeavors.
Acknowledgments. The authors thank Giuseppe Ateniese for the fruitful discussion and insights. This research work was partly funded by MUR under PRIN grant B87G22000450001 (PINPOINT), by the Latium Region under PO FSE+ grant B83C22004050009 (PPMPP), and by the EU-NGEU under the NRRP MUR grant PE00000014 (SERICS).

### 09 Apr 2024 v3

@@ 5.2 The CONFINE protocol
We orchestrate the interaction of the components in CONFINE via a protocol, which consists of four subsequent stages: (i) initialization, (ii) remote attestation, (iii) data transmission, and (iv) computation. These stages are depicted in Figs. 4(a), 4(b), 4(c) and 6, respectively. They are mainly enacted by a Miner Node (multiple instances of which can be deployed in a decentralized fashion) and Provisioner Nodes. We assume their communication channel is reliable [9] and secure [23]. In the following, we describe each of the above phases in detail.
@@
Remote attestation. The remote attestation serves the purpose of establishing trust between miners and provisioners in the context of fulfilling data requests. This phase adheres to the overarching principles outlined in the RATS RFC standard [8] serving as the foundation for several TEE attestation schemes (e.g., Intel EPID111sgx101.gitbook.io/sgx101/sgx-bootstrap/attestation. Accessed: April 9, 2024. and AMD SEV-SNP222amd.com/en/processors/amd-secure-encrypted-virtualization. Accessed: April 9, 2024.). Remote attestation has a dual objective: (i) to furnish provisioners with compelling evidence that the data request for an event log originates from a Trusted App running within a TEE; (ii) to confirm the specific nature of the Trusted App as an authentic Secure Miner software entity. This phase is triggered when the Log Requester sends a new case request to the Log Server, specifying: (i) the segment size (henceforth, seg_size), and (ii) the set of the requested case references. Both parameters will be used in the subsequent data transmission phase. Each of the Log Servers commences the verification process by requesting the necessary information from the Log Receiver to conduct the attestation (2). Subsequently, the Log Receiver generates the attestation report containing the so-called measurement of the Trusted App, which is defined as the hash value of the combination of its source code and data. Once this report is signed using the attestation private key associated with the TEE’s hardware of the Miner Node, it is transmitted by the Log Receiver to the Log Servers alongside the attestation public key of the Miner Node (3). The Log Servers authenticate the miner using the public key and decrypt the report (4). In this last step, the Log Servers undertake a comparison procedure in which they juxtapose the measurement found within the decrypted report against a predefined reference value associated with the source code of the Secure Miner. If the decrypted measurement matches the predefined value, the Miner Node gains trust from the provisioner.
@@
Data transmission. Once the trusted nature of the Trusted App is verified, the Log Servers proceed with the transmission of their cases. To accomplish this, each Log Server retrieves the event log from the PAIS (steps 1 and 2, in Fig. 4(c)), and filters it according to the case reference set specified by the miner. Given the constrained workload capacity of the TEE, Log Servers could be requested to partition the filtered event log into distinct segments. Log segments contain a variable count of entire cases (3). The cumulative size of these segments is governed by the threshold parameter specified by the miner in the initial request (step 1 of the remote attestation phase, Fig. 4(b)). As an illustrative example from our motivating scenario, the Log Server of the hospital may structure the segmentation such that and are in the same segment, whereas the specialized clinic might have and in separate segments. Subsequently, the Log Servers transmit their encrypted segments to the Log Receiver of the Trusted App (4). The Log Receiver, in turn, collects the responses in a queue, processing them one at a time. After decrypting a processed segment (5), the Log Receiver forwards the cases contained therein to the Log Manager (6). Data belonging to the same process instance are merged by the Log Manager to build a single trace (e.g., ) comprehensive of all the events in the partial traces (, and ). To do so, the Log Manager applies a specific merging schema (i.e., a rule specifying the attributes that identify a case) as stated in [10]. In our illustrative scenario, the merging schema to combine the cases of Alice is contingent upon the linkage established through their case identifier (312). We underline that our solution facilitates the incorporation of diverse merging schemas encompassing distinct trace attributes. The outcomes arising from merging the cases within the processed segments are securely stored by the Log Manager in the TEE.
@@
Computation. The Trusted App requires all the provisioners to have delivered data referring to the same process instances. For example, when the hospital and the other organizations have all delivered their information concerning case 312 to the Trusted App, the process instance associated with Alice becomes eligible for the computation phase, illustrated in Fig. 6. The Log Manager forwards the cases earmarked for computation to the Log Elaborator (step 1). These cases may constitute either the entire merged event log or a subset thereof. The former setting entails a single computation routine, thus saving execution time but requiring a larger memory buffer in the TEE, whereas the latter necessitates multiple consecutive elaborations with a lower demand for space. Subsequently, the Log Elaborator proceeds to input the merged cases into the process mining algorithm (2). Notice that the above choice on the buffering of cases affects the selection of the mining algorithm to employ. If we elaborate subsequent batches, each containing a part of all merged cases, the mining algorithm must support incremental processing, enriching the output as new batches come along. An example of this class of algorithms is the HeuristicsMiner [36]. Otherwise, incrementality is not required. Ultimately, the outcome of the computation is relayed by the Log Elaborator from the TEE to the TEE Interface running atop the Operating System of the Miner Node (3). In our motivating scenario, the university and the national institute of statistics, serving as miners, disseminate the outcomes of computations, generating analyses that benefit the provisioners, although the original data are never revealed in clear. Furthermore, our protocol enables the potential for provisioners to have their own Secure Miner, allowing them autonomous control over the computed results. Notice that the CONFINE protocol does not impose restrictions on the post-computational handling of results.
@@ 5.3 Implementation
We implemented the Secure Miner component as an Intel SGX333sgx101.gitbook.io/sgx101/. Accessed: April 9, 2024. trusted application, encoded in Go through the EGo framework.444docs.edgeless.systems/ego. Accessed: April 9, 2024. We resort to a TLS communication channel [34] between miners and provisioners over the HTTP web protocol to secure the information exchange. To demonstrate the effectiveness of our framework, we re-implemented and integrated the HeuristicsMiner discovery algorithm [36] within the Trusted Application. Our implementation of CONFINE, including the HeuristicsMiner implementation in Go, is openly accessible at the following URL: github.com/Process-in-Chains/CONFINE/.
@@ 6 Evaluation
In this section, we evaluate our approach through our implementation. We begin with a convergence analysis to demonstrate the correctness of the data exchange process. As discussed in Sect. 5.1, the availability of space in the dedicated TEE areas is subject to hardware limitations. Therefore, we focus on memory consumption, as exceeding those limits could lower the level of security guaranteed by TEEs. Thus, we gauge the memory usage with synthetic and real-life event logs, to observe the trend during the enactment of our protocol and assess scalability. We discuss our experimental results in the following. All the testbeds and results are available in our public code repository (linked above).
Output convergence. To experimentally validate the correctness of our approach in the transmission and computation phases (see Sect. 5), we run a convergence test. To this end, we created a synthetic event log consisting of cases of events on average (see Table 2) by simulating the inter-organizational process of our motivating scenario (see Fig. 1)555We generated the event log through BIMP (https://bimp.cs.ut.ee/). We filtered the generated log by keeping the sole events that report on the completion of activities, and removing the start and end events of the pharmaceutical company and specialized clinic’s sub-processes. and we partitioned it in three sub-logs (one per involved organization), an excerpt of which is listed in Table 1. We run the stand-alone HeuristicsMiner on the whole log, and processed the sub-logs through our CONFINE toolchain. As expected, the results converge and are depicted in Fig. 7 in the form of a workflow net [1]. For clarity, we have colored activities recorded by the organizations following the scheme of Table 2 (black for the hospital, blue for the pharmaceutical company, and green for the specialized clinic).
@@
Memory usage. Figures 7(a) and 7(b) display plots corresponding to the runtime space utilization of CONFINE (in MegaBytes). Differently from Fig. 7(b), Fig. 7(a) excludes the computation stage by leaving the HeuristicsMiner inactive so as to isolate the execution from the mining-specific operations. The dashed lines mark the starting points for the remote attestation, data transmission and computation stages. We held the segment size (seg_size) constant at MegaBytes. We observe that the data transmission stage reaches the highest peak of memory utilization, which is then partially freed by the subsequent computation stage, steadily occupying memory space at a lower level. To verify whether this phenomenon is due to the synthetic nature of our simulation-based event log, we gauge the runtime memory usage of two public real-world event logs, too: Sepsis [26] and BPIC 2013 [32]. The characteristics of the event logs are summarized in Table 2. Since those are intra-organizational event logs, we split the contents to mimic an inter-organizational context. In particular, we separated the Sepsis log based on the distinction between normal-care and intensive-care paths, as if they were conducted by two distinct organizations. Similarly, we processed the BPIC 2013 log to sort it out into the three departments of the Volvo IT incident management system. Figure 7(c) depicts the results. We observe that the BPIC 2013 log demands the most memory during the initial stages, whereas the Sepsis log is associated with the least expensive run, but the polylines exhibit a matching shape with our synthetic dataset.To verify whether these trends are affected by the dimension of the exchanged data segments, we conducted an additional test to examine memory usage as the seg_size varies. Notably, the polylines displayed in Fig. 7(d) indicate a linear increment of memory occupation until a breakpoint is reached. After that, the memory in use is steady. These points, marked by vertical dashed lines, indicate that the seg_size value that allows the providers to send their whole log partition in a single segment.
@@
Scalability. To examine the scalability of the Secure Miner, we focus on its capacity to efficiently manage an increasing workload in the presence of limited memory resources (as it is the case with TEEs). We set three distinct test configurations by varying our motivating scenario log. In particular, we considered (I) the maximum number of events per case, (II) the number of cases , and (III) the number of provisioning organizations as independent integer variables. To conduct the test on the maximum number of events, we added a loop back from the final to the initial activity of the process model, progressively increasing the number of iterations at a step of , resulting in events. Concerning the test on the number of cases, we simulated additional process instances so that having . Finally, for the assessment of the number of organizations, the test necessitated the distribution of the process model activities’ into a variable number of pools, each representing a different organization (). We parameterized the above configurations with three segment sizes (in KiloBytes): for tests I and II, and for test III (the range is reduced without loss of generality to compensate the partitioning of activities into multiple organizations). To facilitate a more rigorous interpretation of the output trends across varying seg_sizes, we employ two well-known statistical measures. As a primary measure of goodness-of-fit, we employ the coefficient of determination [5], which assesses the degree to which the observed data adheres to the linear () and logarithmic () regressions derived from curve fitting approximations. To delve deeper into the analysis of trends exhibiting a high , we consider the slope of the approximated linear regression [4].
@@
Table 9(d) lists the obtained measurements, which we use to elucidate the observed patterns. Figure 8(a) depicts the results of test I, focusing on the increase of memory utilization when the number of events in the logs grows. We observe that the memory usage trends for seg_size set to and (depicted by green and lilac lines, respectively) are almost superimposable, whereas the setting with (blue line) exhibits significantly higher memory usage. With seg_size assigned with and , approaches , signifying an almost perfect approximation of the linear relation. With these settings, is very low yet higher than , thus indicating that memory usage is likely to continue increasing as the number of maximum events grows. The configuration with yields a higher value, thus suggesting a logarithmic trend, hence a greater likelihood of stabilizing memory usage growth rate as the number of maximum events increases. In Fig. 8(b), we present the results of test II, assessing the impact of the number of cases on memory consumption. As expected, the configurations with seg_size set to and demand lower memory than settings with . The score when seg_size is assigned with and indicate a strong linear relationship between the dependent and independent variables compared to the trend with , which is better described by a logarithmic regression (). Differently from test I, the score associated with the linear approximations with seg_size set to and approaches , indicating that the growth rate of memory usage as the number of cases increases is negligible. In Fig. 8(c), we present the results of test III, on the relation between the number of organizations and memory usage. The chart shows that memory usage trends increase as provisioning organizations increase for all three segment sizes. The values for the three seg_sizes are very high, indicating a strong positive linear correlation. The test with exhibits the slowest growth rate, as corroborated by the lowest (). For the configuration with , the memory usage increases slightly faster (). With , the overall memory usage increases significantly faster than the previous configurations (). We derive from these findings that the Secure Miner may encounter scalability issues when handling settings with a large number of provisioning organizations. Further investigation is warranted to determine the precise cause of this behavior and identify potential mitigation strategies.
In the next section, we discuss other future endeavors stemming from our work.
@@ 7 Conclusion and Future Work
In this paper, we described CONFINE, a decentralized approach to process mining. Based upon TEEs, it guarantees the secrecy and confidentiality of data transmitted to and elaborated by processing nodes outside the perimeter of the event log providers. Our research can spur a number of future investigations and improvements in the field. First, we aim to enhance our solution by readjusting it to the relaxation of underlying assumptions we made, including fair conduct by data provisioners, the absence of injected or maliciously manipulated event logs, the exchange of messages through reliable communication channels where no loss or bit corruption occurs, and the existence of a universal clock for timestamps. Also, we are extending our analysis with formal proofs of soundness and completeness of the protocol. Our future work encompasses the integration of usage control policies that specify rules on event logs’ utilization, too. We plan to design enforcement and monitoring mechanisms to achieve this goal following the principles adopted in [6, 7]. We remark that a possibly severe threat to data secrecy lies in the reconstruction of the original input information back from the mining output. Keeping this aspect in mind is crucial to determine the mining algorithm to be embedded in the Secure Miner. Studies in this regard have been conducted, among others, in [35, 19]. Integrating the proposed recommendations with CONFINE paves the path for future investigations. Finally, we acknowledge that the focus of our implementation is on a specific process discovery task. Nevertheless, our approach has the potential to seamlessly cover a wider array of discovery techniques as well as other process mining functionalities like conformance checking and performance analysis. Showing their integrability with our approach, and drawing guidelines on the use of different algorithms, are research directions we plan to follow.
Acknowledgments. The authors thank Giuseppe Ateniese for the fruitful discussion and insights. This work was partly funded by MUR under PRIN grant B87G22000450001 (PINPOINT), the Latium Region under PO FSE+ grant B83C22004050009 (PPMPP), and Sapienza University of Rome under grant RG123188B3F7414A (ASGARD).